# Cost-aware Stopping for Bayesian Optimization

### v1

@@ Abstract
In automated machine learning, scientific discovery, and other applications of Bayesian optimization, deciding when to stop evaluating expensive black-box functions is an important practical consideration. While several adaptive stopping rules have been proposed, in the cost-aware setting they lack guarantees ensuring they stop before incurring excessive function evaluation costs. We propose a cost-aware stopping rule for Bayesian optimization that adapts to varying evaluation costs and is free of heuristic tuning. Our rule is grounded in a theoretical connection to state-of-the-art cost-aware acquisition functions, namely the Pandora’s Box Gittins Index (PBGI) and log expected improvement per cost. We prove a theoretical guarantee bounding the expected cumulative evaluation cost incurred by our stopping rule when paired with these two acquisition functions. In experiments on synthetic and empirical tasks, including hyperparameter optimization and neural architecture size search, we show that combining our stopping rule with the PBGI acquisition function consistently matches or outperforms other acquisition-function–stopping-rule pairs in terms of cost-adjusted simple regret, a metric capturing trade-offs between solution quality and cumulative evaluation cost.
footnotetext: Code available at: https://github.com/QianJaneXie/CostAwareStoppingBayesOpt.
@@ 1 Introduction
Bayesian optimization is a framework designed to efficiently find approximate solutions to optimization problems involving expensive-to-evaluate black-box functions, where derivatives are unavailable. Such problems arise in applications like hyperparameter tuning [21], robot control optimization [18], and material design [26]. It works by iteratively, (a) forming a probabilistic model of the black-box objective function based on data collected thus far, then (b) optimizing a function called the acquisition function, which balances exploration-exploitation tradeoffs, to carefully choose a new point at which to observe the unknown function in the next iteration.
@@
Several stopping rules have been proposed for Bayesian optimization. Simple heuristics—such as fixing a maximum number of iterations or stopping when improvement falls below a threshold—are widely used in practice [16, 17], but can either stop too early or lead to unnecessary evaluations. Other approaches include acquisition-function-based rules that stop when, for instance, the probability of improvement, expected improvement, or knowledge gradient drop below a preset threshold [16, 19, 8]; confidence-bound-based methods like UCB-LCB [17]; the expected minimum simple regret test [11]; as well as methods based on probabilistic regret bounds (PRB) [24]. Most of these target the classical setting which does not explicitly incorporate function evaluation costs.
In this work, we study how to design cost-aware stopping rules, motivated by two primary factors. First, state-of-the-art cost-aware acquisition functions such as the Pandora’s Box Gittins Index (PBGI) [25] and log expected improvement per cost (LogEIPC) [1] have not yet been studied in the adaptive stopping setting. This is important because—as our experiments in Section 4 will show—for best performance, one should pair different acquisition functions with different stopping rules. Second, while certain stopping rules, such as UCB-LCB, are guaranteed to achieve a low simple regret, they are not necessarily guaranteed to do so with low evaluation costs. This is important because—as our experiments will show—UCB-LCB will often incur high evaluation costs, resulting in a high cost-adjusted simple regret.
@@
    Theoretical Guarantees. We prove in Theorem 1 that our stopping rule, when paired with the PBGI or LogEIPC acquisition functions, satisfies a theoretical upper bound on expected cumulative cost up to stopping, which constitutes the first theoretical guarantee of this type for any adaptive stopping rule for Bayesian optimization.
@@
    Empirical Validation. Our experiments show that combining the PBGI acquisition function with our proposed stopping rule consistently matches or outperforms other combinations of acquisition functions and stopping rules.
@@ 2 Bayesian Optimization and Adaptive Stopping
In black-box optimization, the goal is to find an approximate optimum of an unknown objective function $f:X\to\mathbb{R}$ using a limited number of function evaluations at points $x_{1},\ldots,x_{T}\in X$ where $X$ is the search space and $T$ is a given search budget, potentially chosen adaptively. The convention measures performance in terms of expected simple regret [9, Sec. 10.1], which is given by
@@
where the expectation is taken over all sources of randomness, including the sequence of points $x_{1},\ldots,x_{T}$ selected by the algorithm. Bayesian optimization approaches this problem by building a probabilistic model of $f$—typically a Gaussian process [20], conditioned on the observed data points $(x_{t},y_{t})_{t=1}^{T}$, where $y_{t}=f(x_{t})$. An acquisition function $\alpha_{t}:X\to\mathbb{R}$ then guides the selection of new samples by carefully balancing the exploration-exploitation tradeoff arising from uncertainty about $f$.
@@ 2.1 Cost-aware Bayesian Optimization
Cost-aware Bayesian optimization [13, 2] extends the above setup to account for the fact that evaluation costs can vary across the search space. For instance, in the hyperparameter tuning example of Section 1, costs can vary according to the time needed to train a machine learning model under a given set of hyperparameters $x$. Cost-aware Bayesian optimization handles this by introducing a cost function $c:X\to\mathbb{R}^{+}$, which may be known or unknown depending on the setup. The cost function, or observed costs, are then used to construct the acquisition function $\alpha_{t}$.
In this work, we focus primarily on the cost-per-sample formulation [6, 5, 25] of cost-aware Bayesian optimization, which seeks methods achieving a low expected cost-adjusted simple regret
@@
One can also work with the expected budget-constrained formulation [25], which incorporates budget constraints explicitly, and seeks algorithms which achieve a low simple regret under an expected evaluation budget. Here, performance is evaluated in terms of
$\mathcal{R}$ $=\mathbb{E}\left[\min_{1\leq t\leq T}f(x_{t})-\inf_{x\in X}f(x)\right]$ where $\text{where }x_{1},\ldots,x_{T}\text{ satisfy}$ $\mathbb{E}\sum\limits_{t=1}^{T}c(x_{t})$ $\leq B.$ (3)
The stopping rules we study can be applied in this setting as well: we discuss this in Section 3.2.
For both settings, we work with a number of acquisition functions—chiefly, log expected improvement per cost (LogEIPC) [1] and the Pandora’s Box Gittins Index (PBGI) [25]. These, respectively, are
@@
where $y^{*}_{1:t}=\min_{1\leq s\leq t}y_{s}$ is the best value observed in the first $t$ steps, and $\mathrm{EI}_{\psi}(x;y)=\mathbb{E}\left[\max(y-\psi(x),0)\right]$ is the expected improvement at point $x$ with respect to some random function $\psi:X\to\mathbb{R}$ and a baseline value $y$. In the uniform-cost setting, we also consider the classical lower confidence bound and Thompson sampling acquisition functions; see Garnett [9] for details.
@@ 2.2 Adaptive Stopping Rules
To the best of our knowledge, stopping rules for Bayesian optimization have only been studied in the classical setting without explicit evaluation costs. We broadly categorize existing methods as follows:
@@
    Criteria based on convergence or significance of improvement. This includes empirical convergence, namely stopping when the best value remains unchanged for a fixed number of iterations, or the global stopping strategy (GSS) [3], which stops when the improvement is no longer statistically significant relative to the inter-quartile range (i.e., the range between the 25th percentile and the 75th percentile) of prior observations.
@@
    Acquisition-based criteria. This includes stopping rules built from acquisition functions such as the probability of improvement (PI) [16], expected improvement (EI) [15, 19, 11], and knowledge gradient (KG) [8]. These approaches typically stop when the acquisition value falls below a fixed threshold—a predetermined constant, median of the initial acquisition values, or the cost of sampling—but typically assume uniform costs and do not adapt to evaluation costs which can vary across the search space.
@@
    Regret-based criteria. This includes stopping rules based on confidence bounds such as UCB-LCB [17] and the gap of expected minimum simple regrets [11]. These stop when certain regret bounds fall below a preset or data-driven threshold. The related probabilistic regret bound (PRB) stopping rule [24] stops when a candidate point is estimated to have simple regret below a small threshold $\epsilon$ with confidence $1-\delta$.
In settings beyond Bayesian optimization, Chick and Frazier [6], have proposed a cost-aware stopping rules for finite-domain sequential sampling problems with independent values. Although this formulation allows for varying costs, it does not extend to general Bayesian optimization settings which use correlated Gaussian process models.
@@ 3 A Stopping Rule Based on the Pandora’s Box Gittins Index
In this work, we propose a new stopping rule tailored for two state-of-the-art acquisition functions used in cost-aware Bayesian optimization: LogEIPC and PBGI, introduced in Section 2. As discussed by Xie et al. [25], these acquisition functions are closely connected, arising from two different approximations of the intractable dynamic program which defines the Bayesian-optimal policy for cost-aware Bayesian optimization. We now show that this connection can be used to obtain a principled stopping criterion to be paired with both acquisition functions.
@@
To derive a stopping rule for PBGI, consider first the Pandora’s Box problem, from which it is derived. Pandora’s Box can be seen as a special case of cost-per-sample Bayesian optimization, as introduced in Section 2, where $X=\{1,\dots,N\}$ is a discrete space and $f$ is assumed uncorrelated. In this setting, the observed values do not affect the posterior distribution—meaning, we have $f(x^{\prime})\mid x_{1:t},y_{1:t}=f(x^{\prime})$ at all unobserved points $x^{\prime}$—and thus the acquisition function value at point $x$ is time-invariant and can be written simply as $\alpha^{\mathrm{PBGI}}(x)$, where $\mathrm{EI}_{f}(x;\alpha^{\mathrm{PBGI}}(x))=c(x)$. Following Weitzman [23], one can show using a Gittins index argument that selecting $x_{t}$ to minimize $\alpha^{\mathrm{PBGI}}$ is Bayesian-optimal under minimization—the algorithm which does so achieves the smallest expected cost-adjusted simple regret, among all adaptive algorithms.
@@
One critical detail applied in the optimality argument of Weitzman [23] is that the policy defined by $\alpha^{\mathrm{PBGI}}$ is not Bayesian-optimal for any fixed deterministic $T$. Instead, it is optimal only when the stopping time $T$ is chosen according to the condition [^1] Following Xie et al. [25, Appendix B], optimality holds under any tie-breaking rule in the cost-per-sample setting, but only under a carefully-chosen stochastic tie-breaking rule in the expected budget-constrained setting. For simplicity, we use the stopping-as-early-as-possible tie-breaking rule throughout this paper.
$\min_{x\in X\backslash\{x_{1},\dots,x_{t}\}}\alpha^{\mathrm{PBGI}}(x)\geq y^{*}_{1:t},$ (6)
where, as before, $y^{*}_{1:t}$ is the best value observed so far.
@@
In the correlated setting we study, Xie et al. [25] extend $\alpha^{\mathrm{PBGI}}$ to define an acquisition function by recomputing it at each time step $t$ based on the posterior mean and variance, which defines $\alpha^{\mathrm{PBGI}}_{t}$ as in (5). In order to also extend the associated Pandora’s Box stopping rule, a subtle design choice arises: should we use $\alpha_{t-1}^{\mathrm{PBGI}}$ (before update) or $\alpha_{t}^{\mathrm{PBGI}}$ (after update) in (6)? While prior theoretical work [10] adopts the former choice, we instead propose the latter, for two main reasons.
@@
First, we argue it more faithfully reflects the intuition behind Weitzman’s original stopping rule. In the independent setting, $\alpha^{\mathrm{PBGI}}(x)$ represents a kind of fair value for point $x$ [12]. For evaluated points $x\in\{x_{1},\dots,x_{t}\}$, this is simply the observed function value. For unevaluated points $x\not\in\{x_{1},\dots,x_{t}\}$, the fair value reflects uncertainty in $f(x)$ and the cost $c(x)$. The stopping rule (6) then says to stop when the best fair value is among the already-evaluated points—namely, when no point provides positive expected gain relative to its cost if evaluated in the next round, conditioned on current observations. In the correlated setting, the fair value naturally extends to $\alpha_{t}^{\mathrm{PBGI}}$, because this incorporates all known information at a given time. Second, we show in Section C.2 that using $\alpha_{t}^{\mathrm{PBGI}}$ yields tangible empirical gains in cost-adjusted regret.
@@
We call the stopping rule given by the equivalent conditions (6), (9), and (10) the PBGI/LogEIPC stopping rule. Figure 1 gives an illustration of how the rule behaves in a simple setting, demonstrating that it is more conservative—in the sense of preferring to stop earlier—when the cost is high.
@@
3.1 A Theoretical Guarantee on Cost up to Stopping
In the discrete Pandora’s Box setting, maximizing $\alpha_{t}^{\mathrm{PBGI}}(x)$ is Bayesian-optimal in more than one way: doing so not only achieves the best possible expected cost-adjusted simple regret, under appropriately-scaled costs, it also achieves the best expected simple regret, among all algorithms satisfying an expected budget constraint of the form
$\mathbb{E}\left[\sum_{t=1}^{T}c(x_{t})\right]\leq B,$ (11)
where $B>0$ is the total budget, and the cost function used in defining $\alpha_{t}^{\mathrm{PBGI}}(x)$ is $\lambda c(x)$ for some cost-scaling factor $\lambda$ which depends on $B$. A proof is given by Xie et al. [25, Theorem 2].
This result, at first, appear to be completely Pandora’s-Box-theoretic: it requires $X$ to be discrete and $f$ to be uncorrelated. In the more general correlated setting of cost-aware Bayesian optimization, it is not a priori clear how much costs our stopping rule will incur in expectation. To understand this, we prove the following upper bound on the expected cumulative cost up to the stopping time.

### v2

@@ Abstract
In automated machine learning, scientific discovery, and other applications of Bayesian optimization, deciding when to stop evaluating expensive black-box functions is an important practical consideration. While several adaptive stopping rules have been proposed, in the cost-aware setting they lack guarantees ensuring they stop before incurring excessive function evaluation costs. We propose a cost-aware stopping rule for Bayesian optimization that adapts to varying evaluation costs and is free of heuristic tuning. Our rule is grounded in a theoretical connection to state-of-the-art cost-aware acquisition functions, namely the Pandora’s Box Gittins Index (PBGI) and log expected improvement per cost. We prove a theoretical guarantee bounding the expected cumulative evaluation cost incurred by our stopping rule when paired with these two acquisition functions. In experiments on synthetic and empirical tasks, including hyperparameter optimization and neural architecture size search, we show that combining our stopping rule with the PBGI acquisition function usually matches or outperforms other acquisition-function–stopping-rule pairs in terms of cost-adjusted simple regret, a metric capturing trade-offs between solution quality and cumulative evaluation cost.
@@ 1 Introduction
Bayesian optimization is a framework designed to efficiently find approximate solutions to optimization problems involving expensive-to-evaluate black-box functions, where derivatives are unavailable. Such problems arise in applications like hyperparameter tuning (Snoek et al., 2012), robot control optimization (Martinez-Cantin, 2017), and material design (Zhang et al., 2020). It works by iteratively, (a) forming a probabilistic model of the black-box objective function based on data collected thus far, then (b) optimizing an acquisition function, which balances exploration-exploitation tradeoffs, to carefully choose a new point at which to observe the unknown function in the next iteration.
@@
Several stopping rules have been proposed for Bayesian optimization. Simple heuristics—such as fixing a maximum number of iterations or stopping when the best value remains unchanged for a certain number of iterations—are widely used in practice, but can either stop too early or lead to unnecessary evaluations. Other approaches include acquisition-function-based rules that stop when, for instance, the probability of improvement, expected improvement, or knowledge gradient drop below a preset threshold (Lorenz et al., 2015; Nguyen et al., 2017; Frazier & Powell, 2008); and regret-bound-based rules, which use theoretical performance guarantees to decide termination (Makarova et al., 2022; Ishibashi et al., 2023; Wilson, 2024). Most of these target the classical setting which does not explicitly incorporate function evaluation costs.
In this work, we study how to design cost-aware stopping rules, motivated by two primary factors. First, state-of-the-art cost-aware acquisition functions such as the Pandora’s Box Gittins Index (PBGI) (Xie et al., 2024) and log expected improvement per cost (LogEIPC) (Ament et al., 2023) have not yet been studied in the adaptive stopping setting. This is important because—as our experiments in Section˜4 will show—for best performance, one should pair different acquisition functions with different stopping rules. Second, while certain stopping rules, such as UCB-LCB (Makarova et al., 2022), are guaranteed to achieve a low simple regret, they are not necessarily guaranteed to do so with low evaluation costs. This is important because—as our experiments will show—UCB-LCB will often incur high evaluation costs, resulting in a high cost-adjusted simple regret.
@@
    Theoretical Guarantees. We prove in Theorem˜2 that our stopping rule, when paired with the PBGI or LogEIPC acquisition functions, satisfies a theoretical upper bound on the expected cost-adjusted simple regret, which constitutes the first theoretical guarantee of this type for any adaptive stopping rule for Bayesian optimization.
@@
    Empirical Validation. Our experiments show that combining the PBGI acquisition function with our proposed stopping rule usually matches or outperforms other combinations of acquisition functions and stopping rules.
@@ 2 Bayesian Optimization and Adaptive Stopping
In black-box optimization, the goal is to find an approximate optimum of an unknown objective function $f:X\to\mathbb{R}$ using a limited number of function evaluations at points $x_{1},\ldots,x_{T}\in X$ where $X$ is the search space and $T$ is a given search budget, potentially chosen adaptively. The convention measures performance in terms of expected simple regret (Garnett, 2023, Sec. 10.1), given by
@@
where the expectation is taken over all sources of randomness, including the sequence of points $x_{1},\ldots,x_{T}$ selected by the algorithm. Bayesian optimization approaches this problem by building a probabilistic model of ff—typically a Gaussian process Rasmussen & Williams (2006), conditioned on the observed data points $(x_{t},y_{t})_{t=1}^{T}$, where $y_{t}=f(x_{t})$. An acquisition function $\alpha_{t}:X\to\mathbb{R}$ then guides the selection of new samples by carefully balancing the exploration-exploitation tradeoff arising from uncertainty about $f$.
@@ 2.1 Cost-aware Bayesian Optimization
Cost-aware Bayesian optimization (Lee et al., 2020; Astudillo et al., 2021) extends the above setup to account for the fact that evaluation costs can vary across the search space. For instance, in the hyperparameter tuning example of Section˜1, costs can vary according to the time needed to train a machine learning model under a given set of hyperparameters $x$. Cost-aware Bayesian optimization handles this by introducing a cost function $c:X\to\mathbb{R}^{+}$, which may be known or unknown ahead. The cost function, or observed costs, are then used to construct the acquisition function $\alpha_{t}$.
In this work, we focus primarily on the cost-per-sample formulation (Chick & Frazier, 2012; Cashore et al., 2016; Xie et al., 2024) of cost-aware Bayesian optimization, which seeks methods with stopping time $\tau$, not necessarily fixed, that achieve a low expected cost-adjusted simple regret
@@
One can also work with the expected budget-constrained formulation (Xie et al., 2024), which incorporates budget constraints explicitly, and seeks algorithms which achieve a low simple regret under an expected evaluation budget. Here, performance is evaluated in terms of
$\mathcal{R}$ $=\mathbb{E}\left[\min_{1\leq t\leq{\tau}}f(x_{t})-\inf_{x\in X}f(x)\right]$ $\text{where }x_{1},\ldots,x_{\tau}\text{ satisfy}$ $\mathbb{E}\sum\limits_{t=1}^{\tau}c(x_{t})$ $\leq B.$ (3)
The stopping rules we study can be applied in this setting as well: we discuss this in Section˜3.1.
For both settings, we work with a few acquisition functions—chiefly, log expected improvement per cost (LogEIPC) (Ament et al., 2023) and the Pandora’s Box Gittins Index (PBGI) (Xie et al., 2024):
@@
where $y^{*}_{1:t}=\min_{1\leq s\leq t}y_{s}$ is the best value observed in the first $t$ steps, and $\mathrm{EI}_{\psi}(x;y)=\mathbb{E}\left[\max(y-\psi(x),0)\right]$ is the expected improvement at point $x$ with respect to some random function $\psi:X\to\mathbb{R}$ and a baseline value $y$. In the uniform-cost setting, we also consider the classical lower confidence bound and Thompson sampling acquisition functions; see Garnett (2023) for details.
@@ 2.2 Adaptive Stopping Rules
To the best of our knowledge, stopping rules for Bayesian optimization typically do not incorporate cost explicitly into the stopping criterion. We broadly categorize existing methods as follows:
@@
    Criteria based on convergence or significance of improvement. This includes empirical convergence, namely stopping when the best value remains unchanged for a fixed number of iterations, or the global stopping strategy (GSS) (Bakshy et al., 2018), which stops when the improvement is no longer statistically significant relative to the inter-quartile range (i.e., the range between the 25th percentile and the 75th percentile) of prior observations. In the multi-fidelity setting, Foumani & Bostanabad (2025) proposed stopping when the high-fidelity surrogate’s estimated optimum has stabilized over a window of iterations, which is related to cost-awareness but differs from our setting with explicitly varying evaluation costs.
@@
    Acquisition-based criteria. This includes stopping rules built from acquisition functions such as the probability of improvement (PI) (Lorenz et al., 2015), expected improvement (EI) (Locatelli, 1997; Nguyen et al., 2017; Ishibashi et al., 2023), and knowledge gradient (KG) (Frazier & Powell, 2008). These approaches typically stop when the acquisition value falls below a fixed threshold—a predetermined constant, median of the initial acquisition values, or the cost of sampling—but typically assume uniform costs and do not adapt to evaluation costs which can vary across the search space.
@@
    Regret-based criteria. This includes stopping rules based on confidence bounds such as UCB-LCB (Makarova et al., 2022) and the gap of expected minimum simple regrets (Ishibashi et al., 2023). These stop when certain estimated regret bounds fall below a preset or data-driven threshold. The related probabilistic regret bound (PRB) stopping rule (Wilson, 2024) stops when estimated simple regret is below a small threshold $\epsilon$ with confidence $1-\delta$.
In settings beyond Bayesian optimization, Chick & Frazier (2012) have proposed a cost-aware stopping rules for finite-domain sequential sampling problems with independent values. Although this formulation allows for varying costs, it does not extend to general Bayesian optimization settings which use correlated Gaussian process models.
@@ 3 A Stopping Rule Based on the Pandora’s Box Gittins Index
In this work, we propose a new stopping rule tailored for two state-of-the-art acquisition functions used in cost-aware Bayesian optimization: LogEIPC and PBGI, introduced in Section˜2. As discussed by Xie et al. (2024), these acquisition functions are closely connected, arising from two different approximations of the intractable dynamic program which defines the Bayesian-optimal policy for cost-aware Bayesian optimization. We now show that this connection can be used to obtain a principled stopping criterion to be paired with both acquisition functions.
@@
To derive a stopping rule for PBGI, consider first the Pandora’s Box problem, from which it is derived. Pandora’s Box can be seen as a special case of cost-per-sample Bayesian optimization, as introduced in Section˜2, where $X=\{1,\dots,N\}$ is a discrete space and $f$ is assumed uncorrelated. In this setting, the observed values do not affect the posterior distribution—meaning, we have $f(x^{\prime})\mid x_{1:t},y_{1:t}=f(x^{\prime})$ at all unobserved points $x^{\prime}$—and thus the acquisition function value at point $x$ is time-invariant and can be written simply as $\alpha^{\mathrm{PBGI}}(x)$, where $\mathrm{EI}_{f}(x;\alpha^{\mathrm{PBGI}}(x))=c(x)$. Following Weitzman (1979), one can show using a Gittins index argument that selecting $x_{t}$ to minimize $\alpha^{\mathrm{PBGI}}$ is Bayesian-optimal under minimization—the algorithm which does so achieves the smallest expected cost-adjusted simple regret, among all adaptive algorithms.
@@
One critical detail applied in the optimality argument of Weitzman (1979) is that the policy defined by $\alpha^{\mathrm{PBGI}}$ is not Bayesian-optimal for any fixed deterministic $T$. Instead, it is optimal only when the stopping time $T$ is chosen according to the condition
$min_{x\in X\backslash\{x_{1},\dots,x_{t}\}}\alpha^{\mathrm{PBGI}}(x)\geq y^{*}_{1:t},$ (6)
where $\geq$ can be replaced by >> [^1] Following Xie et al. (2024, Appendix B), optimality holds under any tie-breaking rule in the cost-per-sample setting, but only under a carefully-chosen stochastic tie-breaking rule in the expected budget-constrained setting. For simplicity, we use the stopping-as-early-as-possible tie-breaking rule throughout this paper., and as before, $y^{*}_{1:t}$ is the best value observed so far.
@@
In the correlated setting we study, Xie et al. (2024) extend $\alpha^{\mathrm{PBGI}}$ to define an acquisition function by recomputing it at each time step $t$ based on the posterior mean and variance, which defines $\alpha^{\mathrm{PBGI}}_{t}$ as in (5). In order to also extend the associated stopping rule, a subtle design choice arises: should we use $\alpha_{t-1}^{\mathrm{PBGI}}$ (before update) or $\alpha_{t}^{\mathrm{PBGI}}$ (after update) in (6)? While prior theoretical work (Gergatsouli & Tzamos, 2023) adopts the former choice, we instead propose the latter, for two main reasons.
@@
First, we argue it more faithfully reflects the intuition behind Weitzman’s original stopping rule. In the independent setting, $\alpha^{\mathrm{PBGI}}(x)$ represents a kind of fair value for point $x$ (Kleinberg et al., 2016). For evaluated points $x\in\{x_{1},\dots,x_{t}\}$, this is simply the observed function value. For unevaluated points $x\not\in\{x_{1},\dots,x_{t}\}$, the fair value reflects uncertainty in $f(x)$ and the cost $c(x)$. The stopping rule (6) then says to stop when the best fair value is among the already-evaluated points—namely, when no point provides positive expected gain relative to its cost if evaluated in the next round, conditioned on current observations. In the correlated setting, the fair value naturally extends to $\alpha_{t}^{\mathrm{PBGI}}$, because this incorporates all known information at a given time. Second, we show in Section˜D.2 that using $\alpha_{t}^{\mathrm{PBGI}}$ yields tangible empirical gains in cost-adjusted regret.
@@
We call the stopping rule given by the equivalent conditions (6), (9), and (10) the PBGI/LogEIPC stopping rule. Figure˜1 gives an illustration of how the rule behaves in a simple setting, demonstrating that it is more conservative—preferring to stop earlier—when the cost is high.
@@
3.1 Theoretical Guarantees on Cost-Adjusted Simple Regret
The connection between PBGI and LogEIPC in fact goes beyond a shared stopping rule. In Lemma˜1, we prove that when paired with this stopping rule, both acquisition functions guarantee that at each iteration before stopping, the expected improvement at the selected point is at least its evaluation cost.
Lemma 1.
Let $X$ be compact, and let $f:X\to\mathbb{R}$ be a random function with constant mean $\mu$. Consider a Bayesian optimization algorithm that begins at some initial point $x_{1}\in X$ with cost $C=c(x_{1})$, acquires additional points using either the PBGI or LogEIPC acquisition function, and uses the PBGI/LogEIPC stopping rule. Let $\tau=\min_{t\geq 1}\{\sup_{x\in X}\alpha_{t}^{\mathrm{LogEIPC}}(x)\leq 0\}$ be the algorithm’s stopping time, and denote the posterior expected improvement function by $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$. Then, $\alpha_{t}^{\mathrm{EI}}(x_{t})\geq c(x_{t})$ for all $t\leq\tau$.
Proofs are in Appendix˜B. As a consequence of this claim, we show in Theorem˜2 that the PBGI/LogEIPC stopping rule, when paired with PBGI or LogEIPC, achieve strictly better cost-adjusted simple regret than the naive strategy of stopping immediately after the initial evaluation. The last term in (11) is exactly the expected cost-adjusted simple regret of this stopping-immediately baseline ("Immediate"). While this may sound obvious, many acquisition function–stopping rule pairs fail to satisfy this better-than-immediate property in practice (see Figures˜2, 3 and 4), largely because most stopping rules are designed with simple regret alone in mind and do not explicitly account for evaluation costs. To our knowledge, this is the first result establishing a formal guarantee on cost-adjusted simple regret for Bayesian optimization, as opposed to simpler settings like Pandora’s Box.